# Pólya’s conjecture for thin products

### v2

@@ 3.2. The Dirichlet case with $d_{2}=2$
and the inequality (3.2) gives, for $a^{2}\lambda>C(\Omega_{1})$,
@@
Then by (3.5), (3.6) and (3.7), one has that if $a^{2}\lambda>C(\Omega_{1})$, then
@@
then for any $\lambda>a^{-2}C(\Omega_{1})$, one also gets the demanded inequality
@@
For $\lambda<a^{-2}C(\Omega_{1})$, one just repeat the corresponding part of the proof of the Dirichlet case with $d_{2}\geq 3$, so we omit it. This completes the proof of Theorem 1.1, (1).
@@ 3.3. The Neumann case
It remains to consider the case $a^{-2}\mu_{1}(\Omega_{1})\leq\lambda\leq a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$ assuming $\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))>\mu_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by Corollary 2.4, one has
@@ 3. $\mathrm{(3)}$
Now we start to prove Theorem 1.2. Again by [12, Proposition 3.1], it is enough to assume that $\Omega$ is connected. Since any rectangle in $\mathbb{R}^{2}$ satisfies Pólya’s conjecture, we can assume that the dimension $d$ of $\Omega$ is at least 2. Again we argue by treating $d\geq 3$ and $d=2$ separately, and by treating Dirichlet case and Neumann case separately.
@@ 4.5. The Neumann case with $d\geq 3$
Again by [12, Proposition 3.1], it is enough to assume that both $\Omega$ and $M$ are connected. Let the eigenvalues of $M$ be
@@ 5.3. The Dirichlet case with $d_{1}\geq 2$
If $d_{2}\geq 3$, then as in §3.1, there exists a constant $C(\Omega)>0$ such that for $a^{2}\lambda>C(\Omega)$,
@@
$a<\big{(}\frac{C_{d_{1}+d_{2}-1}|\partial\Omega||M|}{10C_{d_{1}}|\Omega|}\big{)}^{\frac{1}{d_{2}}}\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{-\frac{1}{2}}C(\Omega)^{\frac{d_{2}-1}{2d_{2}}},$
then for any $\lambda>a^{-2}C(\Omega)$, the extra term is controlled by
@@
If $d_{2}=2$, then as in §3.2, there exists a constant $C(\Omega)>0$ such that if $a^{2}\lambda>C(\Omega)$, one has
@@ 5.4. An abstract extension
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with $C^{1}$ boundary and consider the product manifold $a\Omega\times M$.
@@ 6. Two examples with explicit constants
For the triangle $T$, let $P=\{(x,2x)\}_{x\in\mathbb{R}}\cup\{(2x,x)\}_{x\in\mathbb{R}}\cup\{(x,-x)\}_{x\in\mathbb{R}}$, then by [28, Proposition 3], one has
@@
We remark that in [13, Example 2.D], P. Freitas and I. Salavessa had already observed (from a very simple tiling argument) that $(0,a)\times S^{2}$ satisfies Pólya’s inequalities for $a$ small enough but fails to satisfy Pólya’s inequalities for $a$ large. Our method is more complicated but has the advantage that we could give explicit estimates of $a$ for Pólya’s inequalities to hold. We would like to thank the authors for pointing out this fact to us.

### v3

@@ 3.2. The Dirichlet case with $d_{2}=2$
and the inequality (3.2) gives, for $a^{2}\lambda>C_{1}(\Omega_{1})$,
@@
Then by (3.5), (3.6) and (3.7), one has that if $a^{2}\lambda>C_{1}(\Omega_{1})$, then
@@
then for any $\lambda>a^{-2}C_{1}(\Omega_{1})$, one also gets the demanded inequality
@@
For $\lambda<a^{-2}C_{1}(\Omega_{1})$, one just repeat the corresponding part of the proof of the Dirichlet case with $d_{2}\geq 3$, so we omit it.
@@ 3.3. The Neumann case
It remains to consider the case $a^{-2}\mu_{1}(\Omega_{1})\leq\lambda\leq a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$ assuming $\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))>\mu_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by (2.12), one has
@@
Remark 3.2.
As in the Dirichlet case, the dependence of the constant on $\Omega_{1}$ can be made explicit. In fact, if $\Omega_{1}$ is convex, by [31, Theorem 5.3], there exists a constant $C$ depending only on $d_{1}$, such that
(3.13) $\mu_{l}(\Omega_{1})\geq\frac{C}{\operatorname{diam}(\Omega_{1})^{2}}l^{\frac{2}{d_{1}}}$
for all $l$. As a result, if $\Omega_{1}$ is convex, there is no need to introduce $L(\Omega_{1})$ and $C_{2}(\Omega_{1})$ can be selected as a constant depending only on $\operatorname{diam}(\Omega_{1})$ and $d_{1}$. The dependence of $K_{1}(\Omega_{1})$ on $\Omega_{1}$ can be handled as in Remark 3.1. For the dependence of $K_{2}(\Omega_{1})$ on $\Omega_{1}$, one may apply Theorem 1.2 in [12] which indicates that there exists geometric constants $R_{1}$, $R_{2}$ (which depends only on $|\partial\Omega_{1}|$ and $r_{\mathrm{in}}(\Omega_{1})$), such that
$\underset{\mu_{l}(\Omega_{1})<\mu}{\sum}(\mu-\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}\leq L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}-1}{2}}+R_{1}\mu^{\frac{d_{1}+d_{2}-2}{2}}+R_{2}\mu^{\frac{d_{2}-1}{2}}.$
Combine this with (3.13), one gets an explicit constant $K_{2}(\Omega_{1})$ in terms of geometric information of $\Omega_{1}$.
@@ 3. $\mathrm{(3)}$
Now we start to prove Theorem 1.2. Again by [15, Proposition 3.1], it is enough to assume that $\Omega$ is connected. Since any rectangle in $\mathbb{R}^{2}$ satisfies Pólya’s conjecture, we can assume that the dimension $d$ of $\Omega$ is at least 2. Again we argue by treating $d\geq 3$ and $d=2$ separately, and by treating Dirichlet case and Neumann case separately.
@@ 4.5. The Neumann case with $d\geq 3$
Again by [15, Proposition 3.1], it is enough to assume that both $\Omega$ and $M$ are connected. Let the eigenvalues of $M$ be
@@ 5.3. The Dirichlet case with $d_{1}\geq 2$
If $d_{2}\geq 3$, then as in §3.1, there exists a constant $C_{1}(\Omega)>0$ such that for $a^{2}\lambda>C_{1}(\Omega)$,
@@
$a<\big{(}\frac{C_{d_{1}+d_{2}-1}|\partial\Omega||M|}{10C_{d_{1}}|\Omega|}\big{)}^{\frac{1}{d_{2}}}\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{-\frac{1}{2}}C_{1}(\Omega)^{\frac{d_{2}-1}{2d_{2}}},$
then for any $\lambda>a^{-2}C_{1}(\Omega)$, the extra term is controlled by
@@
If $d_{2}=2$, then as in §3.2, there exists a constant $C_{1}(\Omega)>0$ such that if $a^{2}\lambda>C_{1}(\Omega)$, one has
@@ 5.4. An abstract extension
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with Lipschitz boundary and consider the product manifold $a\Omega\times M$.
@@
6. Another class of products satisfying Pólya’s conjecture
As we have mentioned in the introduction, A. Laptev [26] proved that Pólya’s conjecture holds for $\Omega_{1}\times\Omega_{2}$ if it holds for $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ ($d_{2}\geq 2$). In this section, we apply techniques in the proofs of Theorem 1.1 and 1.2 to show that Pólya’s conjecture holds for a class of domains that are close to such products. More precisely, we will show that for any $\Omega_{3}\subset\Omega_{2}$, the difference $\Omega_{1}\times\Omega_{2}-\Omega_{1}\times a\Omega_{3}$ satisfies Pólya’s conjecture for $a$ small enough. The only new input we need is the following well-known fact: If $\Omega_{3}\subset\Omega_{2}$, then
$\mathcal{N}^{D}_{\Omega_{2}\setminus\Omega_{3}}(\lambda)\leq\mathcal{N}^{D}_{\Omega_{2}}(\lambda)-\mathcal{N}^{D}_{\Omega_{3}}(\lambda),$
and if $\Omega_{3}\Subset\Omega_{2}$, then
$\mathcal{N}^{N}_{\Omega_{2}\setminus\Omega_{3}}(\lambda)\geq\mathcal{N}^{N}_{\Omega_{2}}(\lambda)-\mathcal{N}^{N}_{\Omega_{3}}(\lambda).$
Now we state and prove our results. We remark that these theorems also have some abstract version.
Theorem 6.1.
Let $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ be a bounded domain with Lipschitz boundary, $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ ($d_{2}\geq 2$) be a bounded domain which satisfies the Dirichlet Pólya’s conjecture and $\Omega_{3}\subset\Omega_{2}$ be a bounded domain with piece-wise smooth boundary. Then there exists $a_{0}>0$ (depends on $\Omega_{1},\Omega_{2},\Omega_{3}$) such that for any $0<a<a_{0}$, the product $\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})$ satisfies the Dirichlet Pólya’s conjecture (1.2).
Proof.
By (2.6), there exists $C(\Omega_{3})>0$ such that
(6.1) $\mathcal{N}^{D}_{\Omega_{3}}(\lambda)\geq C_{d_{2}}|\Omega_{3}|\lambda^{\frac{d_{2}}{2}}-C(\Omega_{3})\lambda^{\frac{d_{2}-1}{2}},\qquad\forall\lambda>0.$
It follows that for any $\lambda>0$,
$\mathcal{N}^{D}_{(\Omega_{2}\setminus a\Omega_{3})}(\lambda)$ $\leq\mathcal{N}^{D}_{\Omega_{2}}(\lambda)-\mathcal{N}^{D}_{a\Omega_{3}}(\lambda)$
$\leq C_{d_{2}}(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{2}}{2}}+C(\Omega_{3})a^{d_{2}-1}\lambda^{\frac{d_{2}-1}{2}}.$
Now the arguments are similar to those in Section 3.1, 3.2, 4.2 and 4.4. For example, if $d_{1}\geq 2$ and $d_{2}\geq 3$, then as in Section 3.1, there exists $C_{1}(\Omega_{1})>0$ such that
$\mathcal{N}^{D}_{\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})}(\lambda)\leq$ $C_{d_{1}+d_{2}}|\Omega_{1}|(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{1}+d_{2}}{2}}$
$-\frac{1}{5}C_{d_{1}+d_{2}-1}|\partial\Omega_{1}|(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{1}+d_{2}-1}{2}}$
$+L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{3})a^{d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}},\qquad\forall\lambda>C_{1}(\Omega_{1}).$
Thus if we assume
$a<\min\bigg{(}\big{(}\frac{|\Omega_{2}|}{2|\Omega_{3}|}\big{)}^{\frac{1}{d_{2}}},\big{(}\frac{C_{d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{10|\Omega_{1}|C(\Omega_{3})}\big{)}^{\frac{1}{d_{2}-1}}\bigg{)},$
then for any $\lambda>C_{1}(\Omega_{1})$, we will get
$\mathcal{N}^{D}_{\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})}(\lambda)\leq C_{d_{1}+d_{2}}|\Omega_{1}|(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{1}+d_{2}}{2}}.$
We omit the remaining part of the proof. ∎
For Neumann eigenvalues, we have
Theorem 6.2.
Let $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ be a bounded domain with Lipschitz boundary, $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ ($d_{2}\geq 2$) be a bounded domain which satisfies the Neumann Pólya’s conjecture and $0\in\Omega_{3}\Subset\Omega_{2}$ be a bounded domain with piecewise smooth boundary. Then there exists $a_{0}>0$ (depends on $\Omega_{1},\Omega_{2},\Omega_{3}$) such that for any $0<a<a_{0}$, the product $\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})$ satisfies the Neumann Pólya’s conjecture (1.3).
Proof.
Again by (2.6), there exists $C_{1}(\Omega_{3})>0$ such that
(6.2) $\mathcal{N}^{N}_{\Omega_{3}}(\lambda)\leq C_{d_{2}}|\Omega_{3}|\lambda^{\frac{d_{2}}{2}}+C_{1}(\Omega_{3})\lambda^{\frac{d_{2}-1}{2}}+1,\qquad\forall\lambda>0.$
It follows that for any $\lambda>0$,
$\mathcal{N}^{N}_{\Omega_{2}\setminus a\cdot\Omega_{3}}(\lambda)$ $\geq\mathcal{N}^{N}_{\Omega_{2}}(\lambda)-\mathcal{N}^{N}_{a\cdot\Omega_{3}}(\lambda)$
$\geq C_{d_{2}}(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{2}}{2}}-C_{1}(\Omega_{3})a^{d_{2}-1}\lambda^{\frac{d_{2}-1}{2}}-1.$
If $d_{1}\geq 2$, similar to Section 3.3, there exists $C_{1}(\Omega_{1})>0$ such that
$\mathcal{N}^{N}_{\Omega_{1}\times(\Omega_{2}\setminus a\cdot\Omega_{3})}(\lambda)\geq$ $C_{d_{1}+d_{2}}|\Omega_{1}|(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{1}+d_{2}}{2}}$
$+\frac{1}{5}C_{d_{1}+d_{2}-1}(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)|\partial\Omega_{1}|\lambda^{\frac{d_{1}+d_{2}-1}{2}}$
$-C_{1}(\Omega_{3})a^{d_{2}-1}2^{\frac{d_{1}}{2}+1}C_{d_{1}}B(\frac{d_{1}}{2},\frac{d_{2}+1}{2})|\Omega_{1}|d_{1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}$
$-\mathcal{N}^{N}_{\Omega_{1}}(\lambda),\qquad\forall\lambda>C_{1}(\Omega_{1}).$
Next, by (2.5), there exists $C_{2}(\Omega_{1})>0$ such that
$\mathcal{N}^{N}_{\Omega_{1}}(\lambda)\leq 2C_{d_{1}}|\Omega_{1}|\lambda^{\frac{d_{1}}{2}},\qquad\forall\lambda>C_{2}(\Omega_{1}).$
Thus if we assume
$a<\min\bigg{(}\big{(}\frac{|\Omega_{2}|}{2|\Omega_{3}|}\big{)}^{\frac{1}{d_{2}}},\big{(}\frac{C_{d_{1}+d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|}{20C_{1}(\Omega_{3})2^{\frac{d_{1}}{2}+1}C_{d_{1}}B(\frac{d_{1}}{2},\frac{d_{2}+1}{2})|\Omega_{1}|d_{1}}\big{)}^{\frac{1}{d_{2}-1}}\bigg{)},$
then for all
$\lambda>\max\bigg{(}C_{1}(\Omega_{1}),C_{2}(\Omega_{1}),\big{(}\frac{40C_{d_{1}}|\Omega_{1}|}{C_{d_{1}+d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|}\big{)}^{\frac{2}{d_{2}-1}}\bigg{)}=:\Lambda,$
one has
$\mathcal{N}^{N}_{\Omega_{1}\times(\Omega_{2}\setminus a\cdot\Omega_{3})}(\lambda)\geq C_{d_{1}+d_{2}}|\Omega_{1}|(|\Omega_{2}|-a^{d_{2}}|\Omega_{3}|)\lambda^{\frac{d_{1}+d_{2}}{2}}.$
If $\lambda\leq\Lambda$, by (2.12), one has
$\mathcal{N}^{N}_{\Omega_{1}\times\Omega_{2}}(\lambda)$ $=\sum_{\mu_{k}(\Omega_{1})<\lambda}\mathcal{N}^{N}_{\Omega_{2}}(\lambda-\mu_{k}(\Omega_{1}))$
$\geq\sum_{\mu_{k}(\Omega_{1})<\lambda}C_{d_{2}}|\Omega_{2}|(\lambda-\mu_{k}(\Omega_{1}))^{\frac{d_{2}}{2}}$
$>C_{d_{1}+d_{2}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}$
which implies
$\mu_{k}(\Omega_{1}\times\Omega_{2})<\frac{4\pi^{2}}{(\omega_{d_{1}+d_{2}}|\Omega_{1}||\Omega_{2}|)^{\frac{2}{d_{1}+d_{2}}}}k^{\frac{2}{d_{1}+d_{2}}},\qquad\forall k\in\mathbb{Z}_{\geq 0}.$
Since for $a$ small enough, there are at most $\mathcal{N}^{N}_{\Omega_{1}\times\Omega_{2}}(2\Lambda)$ eigenvalues below $\Lambda$, the conclusion follows from the fact that
$\lim_{a\to 0^{+}}\mu_{k}(\Omega_{1}\times(\Omega_{2}\setminus a\cdot\Omega_{3}))=\mu_{k}(\Omega_{1}\times\Omega_{2}),\qquad\forall k\in\mathbb{Z}_{\geq 0}.$
For the case $d_{1}=1$, one just need to modify arguments in Section 4.3 and 4.5 as above. ∎
@@ 7. Two examples with explicit constants
For the triangle $T$, let $P=\{(x,2x)\}_{x\in\mathbb{R}}\cup\{(2x,x)\}_{x\in\mathbb{R}}\cup\{(x,-x)\}_{x\in\mathbb{R}}$, then by [37, Proposition 3], one has
@@
Remark 7.2.
Note that if $\Omega_{a},\Omega_{b}\subset\mathbb{R}^{d}$ which intersect only at boundary, and
$\mathcal{N}^{N}_{\Omega_{a}}(\lambda)\leq C_{d}|\Omega_{a}|\lambda^{\frac{d}{2}}+C_{a}\lambda^{\frac{d-1}{2}},\qquad\mathcal{N}^{N}_{\Omega_{b}}(\lambda)\leq C_{d}|\Omega_{b}|\lambda^{\frac{d}{2}}+C_{b}\lambda^{\frac{d-1}{2}},$
then
$\mathcal{N}^{D}_{\Omega_{a}\cup\Omega_{b}}(\lambda)$ $<\mathcal{N}^{N}_{\Omega_{a}\cup\Omega_{b}}(\lambda)\leq\mathcal{N}^{N}_{\Omega_{a}}(\lambda)+\mathcal{N}^{N}_{\Omega_{b}}(\lambda)$
$\leq C_{d}(|\Omega_{a}|+|\Omega_{b}|)\lambda^{\frac{d}{2}}+(C_{a}+C_{b})\lambda^{\frac{d-1}{2}}.$
In particular, one can calculate $C(\Omega)$ in (2.7) if $\Omega$ is a union of many squares and equilateral triangles, or if $\Omega$ is a union of many $d$-dimensional cubes.
@@
We remark that in [16, Example 2.D], P. Freitas and I. Salavessa had already observed (from a very simple tiling argument) that $(0,a)\times S^{2}$ satisfies Pólya’s inequalities for $a$ small enough but fails to satisfy Pólya’s inequalities for $a$ large. Our method is more complicated but has the advantage that we could give explicit estimates of $a$ for Pólya’s inequalities to hold. We would like to thank the authors for pointing out this fact to us.
Acknowledgment
Funding The authors are partially supported by National Key R and D Program of China 2020YFA0713100, and by NSFC no. 12171446.
Data availability statement This manuscript has no associated data.
Conflict of interest The authors have no Conflict of interest to declare that are relevant to the content of this article.